OpenStudy (anonymous):

Helppp !! Mr. Jones is asked to write two types of radical equations for the SAT Prep Course. The first solution to the radical equation must be extraneous. The second solution to the radical equation must be non-extraneous. Write one equation where the solution is extraneous. Then write a second equation where the solution is non-extraneous. Using complete sentences, explain each step when solving to justify your examples.

OpenStudy (anonymous):

i believe albert einstein is dead if im not mistaken am i right

OpenStudy (anonymous):

Unfortunately, you are correct ...

OpenStudy (anonymous):

what math class are you in

OpenStudy (anonymous):

Algebra 2

OpenStudy (anonymous):

regular not AP not Honors

OpenStudy (anonymous):

Algebra 2 lol ... Regular

OpenStudy (anonymous):

oh man!!!!!! what's wrong with teachers now days they are killing students brain cells

OpenStudy (anonymous):

too much excessive work

OpenStudy (anonymous):

I have little to no brain cells left as a result of mathematics period lol

OpenStudy (anonymous):

LOL well tells mr. Jones that ur contract for Algebra 2 expired LOL im a senior in hig school and i have college readiness Math preparing u for PERT

OpenStudy (anonymous):

Lol i'm a senior also . Well , thanks for lightening the mood a bit but I really gotta get some help to get this done lol .

OpenStudy (anonymous):

well as they always say!!! at first if you don't succeed try and try again!!!!! do i get a medal now? LOL

OpenStudy (anonymous):

Haha you get a medal in my book . But I should probably safe it for someone who answers the question :p

OpenStudy (anonymous):

nice talkin to u though I want to become a fan

OpenStudy (anonymous):

Go for it haha

OpenStudy (anonymous):

will do

OpenStudy (tkhunny):

\(\sqrt{x-4} = -3\)

OpenStudy (anonymous):

@tkhunny i don't know what that is .. you gave me no explanation ..

OpenStudy (tkhunny):

I was just seeing if you were paying attention. What is the first request of the problem statement? "The first solution to the radical equation must be extraneous." That is a quick example of a radical equation with an extraneous solution.

OpenStudy (anonymous):

Ohh ! Okay ! Clever lol .

OpenStudy (anonymous):

@tkhunny ? You left me D:

OpenStudy (kittiwitti1):

.-. I hate story math. --- However, that aside, I must ask you to not explicitly state things like "automatic medal and fan" because that is not the website's purpose. The purpose of OpenStudy is for students to help each other on homework and strengthen their own learning abilities, as well as develop other core skills not commonly practiced in the classroom, NOT helping people for the benefit of fame or followers. Therefore, please remove that part from your post, thanks. I apologize if I sounded rude.

OpenStudy (anonymous):

Lol yeah , that was kinda rude . I have seen that posted everywhere . Nothing gets said about it any other time . There is no rule stating that I cannot say that . Considering they are both FEATURES on this website . I was stating that so everyone was aware .

OpenStudy (tkhunny):

I don't really care about medals or fans. I'm hoping to see you learn the material you need. Can you write a radical equation with both extraneous and non-extraneous solutions?

OpenStudy (anonymous):

I would need to read up on what extraneous means , then I shall be able to answer that for you ! Give me a few moments to make sure I get it right :)

OpenStudy (kittiwitti1):

Well, I did apologize, and I just don't want you to get into trouble, so could you please take it off? :)

OpenStudy (anonymous):

Also , since you're so kind . I will go ahead and award you that fan and medal :) @kittiwitti1 if it helps you sleep at night , sure . I should care about "getting in trouble" , but I don't .. but since you asked so nicely .. why not

OpenStudy (kittiwitti1):

Thanks (: I'm sorry if I offended or annoyed you in any way. I didn't mean it like that :p

OpenStudy (anonymous):

It's fine . Not a big deal really .

OpenStudy (anonymous):

Okay well a radical equation with an extraneous solution could be: 1 = sqrt of x/3 - 2

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Sorry, the website is not working very well. I'm getting no notifications. I jsut have to check back when I can. Is that this \(1 = \sqrt{x/3} - 2\)?

OpenStudy (anonymous):

Omg same here ! And yes , that's correct haha .

OpenStudy (tkhunny):

Let's see... \(3 = \sqrt{x/3}\) 9 = x/3 x = 27 Checking \(3 = \sqrt{27/3}\)?? \(3 = \sqrt{9}\)?? \(3 = 3\)?? That is a non-extraneous solution.

OpenStudy (tkhunny):

If you changed it to +2 you would have a bad one.

OpenStudy (anonymous):

My work is completely different ...

OpenStudy (tkhunny):

Anything like this will do... \(x = \sqrt{something}\) that leads to "Something" less than zero.

OpenStudy (tkhunny):

... or x < 0

OpenStudy (anonymous):

Omg I think I get it now lol . No need to explain further :) I remember doing this now lol . Thank you so much though you really helped a lot !

OpenStudy (tkhunny):

Well, then you are awesome!!! Just a little shove in the right direction. Excellent work.

OpenStudy (anonymous):

Thank you again you have no clue how much I appreciate your help ! I might need your help on one more question though .. Just let me answer this one real quick in my document